# Thermal circuit model for silicon quantum-dot array structures

### v1

@@ I Introduction
In this paper, to clarify the thermal conduction characteristics of QD array structures, we propose a simple thermal circuit model. In Sec. II, we explain our model, where the QD array structure is described by an effective thermal circuit based on the distributed-element model. This results in a analytical expression of the thermal transmission line. In Sec. III, we show the experimental validation of our model, where we measure the electron temperature in a QD array device using CBT. A local heater is implemented by flowing current through the barrier gates of the QD array. This experimental setup mimics the heat generated by the local current flowing through the gate wiring for the qubit addressing Kawakami et al. (2013); Li et al. (2018) and by the microwaves applied to the gate wiring for an electric dipole spin resonance Burkard et al. (2023). From the distance dependence between the local heater and the SET, we find that this model reproduces the experimental results. In Sec. IV, we discuss the limitations of our model and future work. Our proposed model is simple, intuitive, and scalable and can be used for the thermal management of large-scale quantum computers in silicon.
@@ II Model
Figure 1: Conceptual diagram of the heating effects in silicon quantum computers. Heat flow is modeled by describing the electron temperature $T_{e}$ of qubits as a function of the heating power $P$ and various parameters that depend on the device structure.
@@
We describe the heating effects in silicon quantum computers separately in three parts: (a) Heating source, (b) heat flow, and (c) qubits. Figure 1 illustrates an conceptual diagram. In silicon quantum computers that are typically implemented in cryogenic environments using dilution refrigerators, there are several heat sources such as radiation within dilution refrigerator, heat and thermal noise inflows through wires, loss of microwave signals, and local currents in charge sensors Krinner et al. (2019); Savin et al. (2006); Kawakami et al. (2013); Takeda et al. (2018); Noah et al. (2024); Undseth et al. (2023). These heat sources can be modeled as heating power $P$. The heat inflow to the qubits and the heat dissipation to the cooler depend on the device structure, specifically its thermal conductance and capacitance. Our objective is to model the electron temperature of qubits $T_{e}$ as a function of $P$ and various parameters depending on the device structure. By employing this model, we can estimate $T_{e}$ from $P$ to design the heat flow path accordingly.
@@
For constructing the thermal model of the silicon QD array, we note that typical silicon QD array structures are configured with a periodic gate array Vandersypen et al. (2017); Veldhorst et al. (2017); Borsoi et al. (2024); Li et al. (2018); Philips et al. (2022); Lee et al. (2020, 2022), and this periodicity is essential for modeling their thermal characteristics. In our model, the QD array device is represented as a periodic structure shown in Fig. 2(a). The unit cell comprises a set of a plunger gate (PG), barrier gate (BG), and insulators such as SiO2, where the length of the unit cell in $x$ direction is $L_{\rm{cell}}$. For simplicity, the silicon channel and the regions outside the gate structure are omitted. Heat flow is considered two-dimensional due to the presence of thick SiO2 or Si layers at the top and bottom, which have lower thermal conductivity than the gate structures made of metal or polysilicon at cryogenic temperatures. This heat flow is divided into two paths: heat inflow path and dissipation path. The heat inflow path represents the propagation of heat from the source to the qubits along the periodic gate structure, while the dissipation path represents the heat propagation from the heat source to the cooler. Here, the heat flow that does not reach the qubits is dissipated through a medium with high thermal conductivity, such as metallic gates, metallic wires, and metallic structure in a dilution refrigerator.
@@
Figure 2: (a) Model of the silicon QD array structure. The periodic gate structure comprises the plunger gate (PG) and the barrier gate (BG) made of metal or polysilicon. Top and bottom layers are thick Si or SiO2. Heat inflow propagates to the $x$ direction, and heat dissipation propagates to the other directions. (b) Simplified thermal rod model for periodic structure in (a) and the thermal circuit analogous to the transmission line in an electric circuit.
@@
In our model, the thermal structure of the QD array is simplified as a thermal rod, as shown in Fig. 2(b). We employ a thermal circuit model where thermal characteristics are analogously represented to an electrical circuit. In this model, heating power and temperature are represented by current and voltage, respectively, similar to an electrical circuit model Noah et al. (2024). Thus, periodic structures can be modeled using a thermal distributed-element model, analogous to the transmission line model in electrical circuits. We show the thermal circuit diagram for the thermal rod in Fig. 2(b), where unlike in the case of electronic circuits, inductors are excluded. This thermal circuit can be analyzed using the telegrapher’s equations. In this context, $R_{\rm{in}}$ (heat inflow resistance), $R_{\rm{ex}}$ (heat dissipation resistance), and $C_{\rm{ex}}$ (heat capacitance) are defined as distributed elements, as shown in Fig. 2(b).
@@
Here, we outline the assumptions underlying our model. (i) The first assumption is that the thermal circuit consists of an infinite periodic structure. This assumption is valid for large-scale QD array structures and thermally uniform structures such as those incorporating dummy metals in the gate layer. (ii) The second assumption is that the distributed elements of the thermal circuit, $R_{\rm{in}}$, $R_{\rm{ex}}$, and $C_{\rm{ex}}$, are temperature dependent and spatially uniform. To account for the temperature dependence of these spatially uniform circuit elements, we consider the device’s net effective temperature, which is the spatially averaged temperature in the area to be analyzed. (iii) The third assumption is that this net effective temperature is proportional to $T_{e}$. This is an assumption that is also applied in Ref. Noah et al. (2024). These assumptions allow us to use a transmission line model based on the distributed-element model. While the distributed circuit elements are spatially uniform, the spatial temperature distribution of the QD array is not uniform and decays exponentially [as derived in following Eq. (4)]. The approximation error arising from this inconsistency is minimized when the spatial temperature variation is smaller than the net effective temperature.
@@
We then explain the thermal circuit elements $R_{\rm{in}}$, $R_{\rm{ex}}$, and $C_{\rm{ex}}$, respectively. $R_{\rm{in}}$ is the thermal resistance (per unit length) for heat inflow from the heat source to the temperature measurement point (qubit). This thermal resistance is a series of resistances of a periodic structure, which is composed of the metallic or polysilicon gates and the insulator (we show the order estimation of $R_{\rm{in}}$, see Appendix C). On the basis of the above discussion, we use $R_{\rm{in}}\propto 1/T_{e}^{\beta_{\rm{in}}}$, where $\beta_{\rm{in}}$ depends on the thermal conductivity in cryogenic regime and the thickness of the gate and insulator. For example, $\beta_{\rm{in}}\approx 1$ when the gate is dominant and $\beta_{\rm{in}}\approx 2$ when the insulator (amorphous SiO2) is dominant (see Appendix A).
@@
$R_{\rm{ex}}$ is the thermal resistance (inverse of the thermal conductance per unit length) for heat dissipation from the heat source. Heat is dissipated through metallic wires and metallic structures in a dilution refrigerator and is finally collected in the mixing chamber, which acts as a cooler. Therefore, $R_{\rm{ex}}$ depends on the structure of the whole chip and chip’s implementation form such as the circuit board mounting the chip, bonding wires, sample packaging, and cables. Most of these materials are metallic. We then use $R_{\rm{ex}}\propto 1/T_{e}^{\beta_{\rm{ex}}}$, where $\beta_{\rm{ex}}\approx 1$ considering the metallic materials are dominant (see Appendix A).
The heat capacitance $C_{\rm{ex}}$ is crucial for real-time thermal management because it influences the temporal changes in device temperature, which is essential for understanding the dynamic temperature effects on qubits, such as heat-induced frequency shifts Takeda et al. (2018); Undseth et al. (2023); Sato and Kawahara (2024). It is noted that specific heat changes drastically (approximately $\propto T^{3}$) at cryogenic temperatures, according to the Debye model, as shown in Appendix A.
In analogy to transmission line models in electrical circuits, $T_{e}$, the temperature of the thermal rod at a point that is a distance $D$ away from the heat source having heating power $P$, can be derived as follows. From the telegrapher’s equations, we obtain the relations ${T_{e}(x+D)=T_{e}(x)e^{-\gamma_{\rm{th}}D}}$, ${P(x+D)=P(x)e^{-\gamma_{\rm{th}}D}}$, and ${T_{e}(x)=Z_{\rm{th}}P(x)}$, where the position on the thermal rod $x$, thermal characteristic impedance $Z_{\rm{th}}$ and thermal propagation constant $\gamma_{\rm{th}}$ are introduced. From these relations, we can describe
$T_{e}(x+D)=Z_{\rm{th}}e^{-\gamma_{\rm{th}}D}P(x).$ (1)
Since $x$ is arbitrary, we will denote $T_{e}(x+D)$ simply as $T_{e}$ and $P(x)$ as $P$ from here on.
We then describe $Z_{\rm{th}}$ and $\gamma_{\rm{th}}$ as
$Z_{\rm{th}}$ $=\sqrt{\frac{R_{\rm{in}}}{\left(\frac{1}{R_{\rm{ex}}}+j\omega C_{\rm{ex}}\right)}}\approx\sqrt{R_{\rm{in}}R_{\rm{ex}}}\approx a^{1+\beta}/T_{e}^{\beta},$ (2)
$\gamma_{\rm{th}}$ $=\sqrt{R_{\rm{in}}\left(\frac{1}{R_{\rm{ex}}}+j\omega C_{\rm{ex}}\right)}\approx\sqrt{\frac{R_{\rm{in}}}{R_{\rm{ex}}}},$ (3)
where $j$ and $\omega$ are the imaginary unit and angular frequency of thermal behaviors, respectively. The terms $a$ and $\beta$ are constant parameters depending on the device, where $\beta=(\beta_{\rm{in}}+\beta_{\rm{ex}})/2$. In the first approximation in Eqs. (2) and (3), $\omega$ is set to zero to perform the static analysis. The second approximation in Eq. (2) is derived by the relations $R_{\rm{in}}\propto 1/T_{e}^{\beta_{\rm{in}}}$ and $R_{\rm{ex}}\propto 1/T_{e}^{\beta_{\rm{ex}}}$.
Substituting Eq. (2) into Eq. (1), we obtain
$T_{e}\approx ae^{-\frac{\gamma_{\rm{th}}}{1+\beta}D}P^{\frac{1}{1+\beta}}.$ (4)
@@
This result indicates that the electron temperature is modeled as a function of heating power $P$, the distance $D$, and three device-dependent parameters ($a$, $\gamma_{\rm{th}}$, and $\beta$).
Hereafter, for simplicity, we use $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\approx 1$, i.e., $\beta\approx 1$. This approximation is valid because, in a typical QD array structure, the thickness of the insulator is relatively lower than that of the gate, justifying $\beta_{\rm{in}}\approx 1$. Similarly, $\beta_{\rm{ex}}\approx 1$ is valid considering that metallic materials dominate the heat dissipation path, as mentioned above. Therefore, we can rewrite Eq. (4) as
$T_{e}\approx ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P}.$ (5)
@@
We note that $a$ depends on the multiple of the thermal resistances ($R_{\rm{in}}\times R_{\rm{ex}}$) from Eq. (2). On the other hand, $\gamma_{\rm{th}}$ depends on the ratio of the thermal resistances ($R_{\rm{in}}/R_{\rm{ex}}$) from Eq. (3). Furthermore, both $a$ and $\gamma_{\rm{th}}$ are independent of $T_{e}$. The term $2/\gamma_{\rm{th}}$ represents the characteristic length of the heat dissipation, indicating that the heating effect decreases by $1/e$. Thereby, the temperature independent parameters $a$ and $\gamma_{\rm{th}}$ thermally characterize the QD array structure.
Here, we extend our model to the case of multiple heat sources. Typically, the electron temperature is written as $T_{e}^{\prime}=\sqrt{T_{\rm{MXC}}^{2}+T_{e}^{2}}$, where $T_{e}^{\prime}$, $T_{\rm{MXC}}$, and $T_{e}$ are the measurable temperature, the base temperature of refrigerator, and electron temperature, respectively Noah et al. (2024); de Kruijf et al. (2024). Consistent with this notation, we employ the following model, i.e., the total rise in electron temperature $T_{e}^{\rm{total}}$ due to multiple heat sources is calculated using the root sum square as
$T_{e}^{\rm{total}}=\sqrt{\sum_{i=1}^{N}(\Delta T_{e}^{i})^{2}},$ (6)
where $N$ is the number of heat sources and $\Delta T_{e}^{i}$ is the rise in electron temperature due to the $i$-th heat source. From Eq. (6), Eq. (5) can be rewritten under the background temperature $T_{0}$ as
$T_{e}^{\rm{total}}=\sqrt{T_{e}^{2}+T_{0}^{2}}=ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P+P_{0}},$ (7)
@@
where $T_{0}=ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P_{0}}$, and $P_{0}$ is the background heating power. We use this equation to fit experimental results. Note that in our experiment, $T_{0}$ and $P_{0}$ includes the effect of lifetime broadening and charge noise broadening de Kruijf et al. (2024).

### v2

@@ I Introduction
In this paper, to clarify the thermal conduction characteristics of QD array structures, we propose a simple thermal circuit model. In Sec. II, we explain our model, where the QD array structure is described by an effective thermal circuit based on the distributed-element model. This results in an analytical expression of the thermal transmission line. In Sec. III, we show the experimental validation of our model, where we measure the electron temperature in a QD array device using CBT. A local heater is implemented by flowing current through the barrier gates of the QD array. This experimental setup mimics the heat generated by the local current flowing through the gate wiring for the qubit addressing Kawakami et al. (2013); Li et al. (2018) and by the microwaves applied to the gate wiring for an electric dipole spin resonance Burkard et al. (2023). From the distance dependence between the local heater and the SET, we find that this model reproduces the experimental results. In Sec. IV, we discuss the limitations of our model and future work. Our proposed model is simple, intuitive, and scalable and can be used for the thermal management of large-scale quantum computers in silicon.
@@ II Model
Figure 1: Conceptual diagram of the heating effects in silicon quantum computers. Heat flow is modeled by describing the electron temperature $T_{e}$ of qubits as a function of the heating power $P_{0}$ and various parameters that depend on the device structure.
@@
We describe the heating effects in silicon quantum computers separately in three parts: (a) Heat source, (b) heat flow, and (c) qubits. Figure 1 illustrates an conceptual diagram. In silicon quantum computers that are typically implemented in cryogenic environments using dilution refrigerators, there are several heat sources such as radiation within dilution refrigerator, heat and thermal noise inflows through wires, loss of microwave signals, and local currents in charge sensors Krinner et al. (2019); Savin et al. (2006); Kawakami et al. (2013); Takeda et al. (2018); Noah et al. (2024); Undseth et al. (2023). These heat sources can be modeled as heating power $P_{0}$. The heat inflow to the qubits and the heat dissipation to the cooler depend on the device structure, specifically its thermal conductance and capacitance. Our objective is to model the electron temperature of qubits $T_{e}$ as a function of $P_{0}$ and various parameters depending on the device structure. By employing this model, we can estimate $T_{e}$ from $P_{0}$ to design the heat flow path accordingly.
@@
For constructing the thermal model of the silicon QD array, we note that typical silicon QD array structures are configured with a periodic gate array Vandersypen et al. (2017); Veldhorst et al. (2017); Borsoi et al. (2024); Li et al. (2018); Philips et al. (2022); Lee et al. (2020, 2022), and this periodicity is essential for modeling their thermal characteristics. In our model, the QD array device is represented as a periodic structure shown in Fig. 2(a). The unit cell comprises a set of a plunger gate (PG), barrier gate (BG), and insulators such as SiO2, where the length of the unit cell in $x$ direction is $L_{\rm{cell}}$. For simplicity, the silicon channel and the regions outside the gate structure are omitted. Heat is generated at a gate, and a portion of it propagates toward a qubit located some distance away. All of the heat is eventually dissipated outside the chip. Accordingly, this heat flow is divided into two paths: heat inflow path and dissipation path. The heat inflow path represents the propagation of heat from the source to the qubits along the periodic gate structure, while the dissipation path represents the heat propagation from the heat source to the cooler. Here, the heat is dissipated mainly through a medium with relatively high thermal conductivity, such as metallic or polysilicon gates, thick SiO2 or Si at the bottom layer, metallic wires, and metallic structure in a dilution refrigerator.
@@
Figure 2: (a) Model of the silicon QD array structure. The periodic gate structure comprises the plunger gate (PG) and the barrier gate (BG) made of metal or polysilicon. Top and bottom layers are thick Si or SiO2. Heat inflow propagates to the $x$ direction, and heat dissipation propagates to the other directions. This conceptual diagram shows a case where the heat source gate and the measurement point are located at a certain distance from each other. (b) Thermal circuit model from the MXC to the chip in the dilution refrigerator. (c) Thermal circuit model for the periodic silicon QD array structure.
We employ a thermal circuit model where thermal characteristics are analogously represented to an electrical circuit. In this model, heating power and temperature are represented by current and voltage, respectively, similar to an electrical circuit model Noah et al. (2024). The overall dilution refrigerator is modeled to set the base temperature of the QD array device, denoted by $T_{\rm{base}}$, as shown in Fig. 2(b), based on Ref. Noah et al. (2024). Assume that the mixing chamber (MXC) is cooled to a stable temperature $T_{m}$, where the cooling power and heating power are balanced. The effective temperature-dependent thermal resistance, denoted by $R_{m}$, includes the thermal resistance of the MXC plate, the cold finger, the printed circuit board (PCB), and wiring from the MXC plate to the chip.
@@
The periodic structures of the QD array can be modeled using a thermal distributed-element model, analogous to the transmission line model in electrical circuits. We show the thermal circuit diagram for the QD array structure in Fig. 2(c), where unlike in the case of electronic circuits, capacitors and inductors are excluded. This thermal circuit can be analyzed using simultaneous differential equations analogous to the telegrapher’s equations. In this context, $R_{\rm{in}}$ (heat inflow resistance) and $R_{\rm{ex}}$ (heat dissipation resistance) are defined as distributed elements, as shown in Fig. 2(c). The low side reference of this thermal transmission line is $T_{\rm{base}}$ and the electron temperature rise from $T_{\rm{base}}$ at position $x$ is denoted by $T_{e}(x)$.
@@
Here, we outline the assumptions underlying our model. (i) The first assumption is that the thermal circuit consists of a semi-infinite periodic structure. This assumption is valid for large-scale QD array structures and thermally uniform structures such as those incorporating dummy metals in the gate layer. (ii) The second assumption is that the distributed elements of the thermal circuit, $R_{\rm{in}}$ and $R_{\rm{ex}}$ are temperature dependent and spatially varied. To account for the temperature dependence of these spatially non-uniform circuit elements, we consider the device’s local effective temperature. (iii) The third assumption is that this local effective temperature is proportional to $T_{e}(x)$. These assumptions allow us to use a modified transmission line model based on the distributed-element model.
@@
We then explain the thermal circuit elements $R_{\rm{in}}$ and $R_{\rm{ex}}$, respectively. $R_{\rm{in}}$ is the thermal resistance (per unit length) for heat inflow from the heat source to the temperature measurement point (qubit). This thermal resistance is a series of resistances of a periodic structure, which is composed of the metallic or polysilicon gates and the insulator (we show the order estimation of $R_{\rm{in}}$, see Supplementary Material II). On the basis of the above discussion, we can write $R_{\rm{in}}(x)=a_{\rm{in}}/(T_{e}(x)+T_{\rm{base}})^{\beta_{\rm{in}}}$, where $a_{\rm{in}}$ is a constant and $\beta_{\rm{in}}$ depends on the thermal conductivity in cryogenic regime and the thickness of the gate and insulator. Namely, $R_{\rm{in}}(x)$ depends only on $T_{e}(x)$ due to the local effect. For example, $\beta_{\rm{in}}\approx 1$ when the gate is dominant and $\beta_{\rm{in}}\approx 2$ when the insulator (amorphous SiO2) is dominant (see Supplementary Material I and Refs. Duthil (2014); Ekin (2006); ais ).
@@
$R_{\rm{ex}}$ is the thermal resistance (inverse of the thermal conductance per unit length) for heat dissipation from the heat source. Heat is dissipated through metallic wires and metallic structures in a dilution refrigerator and is finally collected in the MXC, which acts as a cooler. Therefore, potentially, $R_{\rm{ex}}$ depends on the structure of the whole chip and chip’s implementation form such as the circuit board mounting the chip, bonding wires, sample packaging, and cables. However, we assume that the temperature dependence of this resistance is mainly determined by the local effective temperature, as mentioned in the assumption (ii). This is because, along the heat propagation path, regions near the gate structure have high thermal resistance, while regions farther away—such as metal wiring layers, the bottom layer, and metallic structures—have relatively low thermal resistance. As a result, large thermal gradients are likely to form near the gate structure, whereas the more distant regions remain close to $T_{\rm{base}}$. This assumption holds when a good heat dissipation structure is present around the gate region. We then write $R_{\rm{ex}}(x)=a_{\rm{ex}}/(T_{e}(x)+T_{\rm{base}})^{\beta_{\rm{ex}}}$, where $\beta_{\rm{ex}}=$1–2, considering the composition of the local region.
We formulate the relations between the electron temperature $T_{e}(x)$ and the heat flow $P(x)$ as following simultaneous differential equations:
$-\frac{dT_{e}(x)}{dx}$ $=R_{\rm{in}}(x)P(x),$ (1)
$-\frac{dP(x)}{dx}$ $=\frac{1}{R_{\rm{ex}}(x)}T_{e}(x).$ (2)
Here, we assume that $T_{e}(x)\gg T_{\rm{base}}$ and $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\,(\equiv\beta)$, which are referred to as assumptions (iv) and (v), respectively. Assumption (iv) is valid when thermal effects are significant—precisely the regime in which this analysis becomes relevant. The validity of assumption (v) is supported when the temperature dependence of the gate and insulator is similar, or when one of the ratios dominates, both in the local region (as considered in assumption (ii)). On the basis of above discussion, we can reasonably set $\beta\approx 1\text{--}2$.
Based on these assumptions, we can rewrite the equations into an analytically solvable form as follows:
$-\frac{dT_{e}(x)}{dx}$ $=\frac{a_{\rm{in}}}{\left[T_{e}(x)\right]^{\beta}}P(x),$ (3)
$-\frac{dP(x)}{dx}$ $=\frac{\left[T_{e}(x)\right]^{\beta+1}}{a_{\rm{ex}}}.$ (4)
Considering boundary conditions of $P(x=0)=P_{0}\geq 0$ and $P(x=\infty)=T_{e}(x=\infty)=0$, we can obtain one of the solutions for $x\geq 0$ and $P(x)\geq 0$ as
$T_{e}(x)$ $=\left[a_{\rm{in}}a_{\rm{ex}}(\beta+1)\right]^{\frac{1}{2(\beta+1)}}e^{-\sqrt{\frac{a_{\rm{in}}}{a_{\rm{ex}}(\beta+1)}}x}P_{0}^{\frac{1}{\beta+1}}$
$=ae^{-\frac{x}{L_{\rm{th}}}}P_{0}^{\frac{1}{\beta+1}},$ (5)
$P(x)$ $=P_{0}e^{-\sqrt{\frac{a_{\rm{in}}(\beta+1)}{a_{\rm{ex}}}}x}.$ (6)
@@
where a coefficient $a\equiv[a_{\rm{in}}a_{\rm{ex}}(\beta+1)]^{\frac{1}{2(\beta+1)}}$ and a thermal characteristic length $L_{\rm{th}}\equiv\sqrt{\frac{a_{\rm{ex}}(\beta+1)}{a_{\rm{in}}}}$ are introduced. The main result in this section, Eq. (5), indicates that the electron temperature is modeled as a function of heating power $P_{0}$, the distance $x$, and three device-dependent parameters ($a$, $L_{\rm{th}}$, and $\beta$).
@@
We note that the parameter $a$ depends on the multiple of the thermal resistances coefficients ($a_{\rm{in}}\times a_{\rm{ex}}$). On the other hand, $L_{\rm{th}}$ depends on the ratio of the thermal resistances coefficients ($a_{\rm{ex}}/a_{\rm{in}}$). Furthermore, both $a$ and $L_{\rm{th}}$ are independent of $T_{e}$. The term $L_{\rm{th}}$ represents the characteristic length of the heat dissipation, indicating that the heating effect decreases by $1/e$. Thereby, the temperature independent parameters $a$ and $L_{\rm{th}}$ thermally characterize the QD array structure.
Here, background heating power is introduced to fit the experimental results using the proposed model, Eq. (5). The total electron temperature rise, denoted by $T_{e}^{\rm{total}}$, is described as
$T_{e}^{\rm{total}}=\left(T_{e}^{\beta+1}+T_{B}^{\beta+1}\right)^{\frac{1}{\beta+1}}=ae^{-\frac{x}{L_{\rm{th}}}}\left(P_{0}+P_{B}\right)^{\frac{1}{\beta+1}},$} (7)
@@
where $T_{B}=ae^{-\frac{x}{L_{\rm{th}}}}P_{B}^{\frac{1}{\beta+1}}$, and $P_{B}$ represents the background heating power. We use this equation to fit the experimental data. Note that in our experiment, $T_{B}$ and $P_{B}$ include the effects of lifetime broadening and charge noise broadening de Kruijf et al. (2024), whereas $P_{0}\gg P_{B}$ typically holds in most cases.
Finally, the validity of the model is confirmed for a finite and discrete structure. Although assumption (i) considers a semi-infinite structure, the proposed model, Eq. (5), also accommodates similar circuit configurations for finite systems. The detailed calculation method is provided in Supplementary Material III. These results indicate that the device structure used in the experiment remains valid, even if it deviates from the semi-infinite configuration assumed in assumption (i).